OpenStudy (chris215):

find the domain of the given function. f(x)=sqrt(x+3)/(x+8)(x-2)

OpenStudy (anonymous):

is all that mess under the radical?

OpenStudy (chris215):

\[f(x)=\sqrt{x+3}/(x+8)(x-2)\]

OpenStudy (anonymous):

only the numerator is under the radical?

OpenStudy (anonymous):

do u go to connecitons 8grade

OpenStudy (chris215):

yes @satellite73

OpenStudy (anonymous):

no i am in ninth grade

OpenStudy (chris215):

and no @ddds

OpenStudy (anonymous):

ok so first lets take care of the numerator since you cannot take the square root of a negative number, solve \[x+3\geq 0\] in one step

OpenStudy (anonymous):

you got that?

OpenStudy (chris215):

x\[x \ge -3\]

OpenStudy (anonymous):

yeah so that is a start

OpenStudy (anonymous):

then set the denominator equal to zero and solve, because you cannot divide by zero i.e solve \[(x+8)(x-2)=0\] what do you get?

OpenStudy (chris215):

x=2,-8?

OpenStudy (anonymous):

right so those values are out but you can ignore the \(-8\) because we already know \(x\) has to be greater than or equal to \(-3\) now we also know it cannot be 2

OpenStudy (anonymous):

then i guess the question is, how do you write that? are you supposed to answer in interval notation? in words? as an inequality?

OpenStudy (chris215):

those are my choices

OpenStudy (chris215):

so would it be the second to last?

OpenStudy (anonymous):

yeah it is always C

OpenStudy (chris215):

ok thanks so much :)

OpenStudy (anonymous):

yw